### v1

# Fast LiDAR Data Generation with Rectified Flows

Abstract
Building LiDAR generative models holds promise as powerful data priors for restoration, scene manipulation, and scalable simulation in autonomous mobile robots. In recent years, approaches using diffusion models have emerged, significantly improving training stability and generation quality. Despite the success of diffusion models, generating high-quality samples requires numerous iterations of running neural networks, and the increasing computational cost can pose a barrier to robotics applications. To address this challenge, this paper presents R2Flow, a fast and high-fidelity generative model for LiDAR data. Our method is based on rectified flows that learn straight trajectories, simulating data generation with much fewer sampling steps against diffusion models. We also propose a efficient Transformer-based model architecture for processing the image representation of LiDAR range and reflectance measurements. Our experiments on the unconditional generation of the KITTI-360 dataset demonstrate the effectiveness of our approach in terms of both efficiency and quality.
I Introduction
LiDAR sensors provide accurate 3D point clouds of surrounding environments by performing an omnidirectional time-of-flight (ToF) ranging. The LiDAR point clouds play an important role for autonomous mobile robots to understand their surroundings both geometrically and semantically, through techniques such as SLAM, object detection, and semantic segmentation. However, the performance of these techniques can be degraded due to incompleteness in adverse weather conditions and point density gaps between different LiDAR sensors. Restoring the degraded point clouds requires a data prior that models complex real-world patterns.
Generative modeling of LiDAR data [1, 2, 3, 4, 5, 6, 7, 8] has been studied for the purposes, motivated by the significant progress in deep generative models [9]. Deep generative models aim to build neural networks that represent the probability density distribution underlying the given samples, e.g., large-scale datasets of LiDAR point clouds. So far, the deep generative models trained on LiDAR data have proven useful as prior knowledge in translation tasks such as sparse-to-dense completion [2, 3, 5, 4] and simulation-to-real (sim2real) domain transfer [3], which also boost the perception tasks such as semantic segmentation. In particular, among the generative model frameworks, diffusion models have led to substantial advancements in LiDAR domain [5, 4, 6, 8] due to their training stability and ability to generate high-quality samples.
Despite the success of diffusion models, generating high-quality LiDAR data incurs significant computational cost. In general, sampling in diffusion models is formulated as a stochastic differential equation (SDE) [10] that describe trajectories from the latent distribution to the data distribution. Accurately simulating these learned trajectories requires hundreds or even thousands of discretized steps, each involving the execution of a deep neural network. Fig. 1 illustrates the trade-offs between the number of sampling steps and the sample quality for recent methods on LiDAR data generation [6, 4]. As shown in Fig. 1, naively reducing the number of sampling steps degrades the quality of the generated LiDAR data. This limitation poses a challenge for robotics applications, where power efficiency and computational speed are critical constraints.
(a) LiDM [6] + APE (b) R2DM [4] (c) R2Flow (ours)
Refer to caption Refer to caption Refer to caption
Feature-space iteration Pixel-space iteration Pixel-space iteration
273M params 31M params 81M params
Refer to captionblurry Refer to caption Refer to caption
NFE = 200 NFE = 256 NFE = 256
(3.7 s) (3.7 s) (7.6 s)
Refer to captioncorrupted Refer to captioncorrupted Refer to caption
NFE = 2 NFE = 2 NFE = 2
(38 ms) (56 ms) (62 ms)
Figure 1: Comparison of range image-based LiDAR generative models. Each column shows a schematic diagram of an iterative generation process (top) and LiDAR point clouds (middle & bottom) generated with the different NFEs (number of function evaluations). In comparison to the baselines (a & b), our proposed R2Flow (c) can generate the consistent and high-quality samples even with the low NFE setting.
To this end, we propose R2Flow (Range–Reflectance Flow), a novel generative model for fast and realistic LiDAR data generation. As a framework for building generative models, we employ rectified flows [11, 12], a kind of conditional flow matching framework for training continuous normalizing flows [13, 14], which have demonstrated generating natural images such as human faces and common objects. Similar to diffusion models, the rectified flows represent the data generation as an iterative process using a deep neural network. Its most distinctive feature is that the process is formulated as deterministic straight trajectories, while the diffusion models employ stochastic curved trajectories. Therefore, the sampling is robust to the number of steps or the step size, and ideally, the entire generation trajectory can be simulated in a single step. Similar to relevant studies [1, 2, 3, 4, 5, 6], R2Flow is trained on the equirectangular image representation of multimodal measurements: range and reflectance (intensity of laser reflection). Moreover, we also propose a neural network architecture to generate the multimodal images based on lightweight vision transformers (ViTs) [15]. Among recent architectures, we verify that our approach achieve the better sample quality for the LiDAR data generation task, suppressing the computational cost and the model size. We conduct an unconditional generation task with the KITTI-360 dataset [16] to evaluate quality and efficiency. Our approach surpasses the state-of-the-art results for both large and small number of steps. We summarize our contributions as follows:
  • We propose R2Flow, a rectified flow-based deep generative model for fast and realistic generation of LiDAR range and reflectance modalities.
  • We also propose a model architecture built upon the recent vision transformers, which is compatible of fidelity and efficiency for LiDAR data generation.
  • The effectiveness of our approach is demonstrated through an unconditional generation task on the KITTI-360 dataset.
II Related Work
Here, we briefly summarize three classes of the existing LiDAR generative models using neural networks.
Variational autoencoders (VAEs). VAEs [17] are trained with an autoencoder with latent representation at the bottleneck. Caccia et al. [1] initiated early work on generating LiDAR range images using a vanilla VAE. In spite of the ELBO-based stable training, the generated samples were blurry. More recently, Xiong et al. [7] employed the improved framework, vector-quantized variational autoencoder (VQ-VAE) [18], with the voxel representation.
Generative adversarial networks (GANs). GANs [19] consist of two competing neural networks, a generator and a discriminator, and have been actively applied in various domains over the last decade. Caccia et al. [1] reported the first results by training a basic GAN on range images. DUSty [2] and DUSty v2 [3] proposed architectures robust to raydrop noise (missing points caused by non-returned laser signals). GANs achieved better quality than VAEs; however, training is unstable, and the resultant point clouds still deviate significantly from real samples.
Diffusion models. In recent years, diffusion models have gained significant attention for their stable training and high-quality sample generation. Diffusion models define the data and latent spaces with the same dimensionality and construct a Markov chain between them over multiple timesteps. Moreover, it is known that the family of diffusion models can be generalized as stochastic differential equations (SDEs) [10]. Here we introduce some studies that proposed the diffusion models for LiDAR data generation. LiDARGen [5] trained range and reflectance images using a variance-exploding SDE [20] trained on pixel-space with a discrete-time schedule. R2DM [4] also trained range and reflectance images but utilized a variance-preserving SDE [21] trained on pixel-space with a continuous-time schedule. The pixel-space diffusion models can capture fine details while they require high computational cost due to the iterative nature of sampling. To mitigate this issue, Ran et al. [6] proposed LiDM with architectural improvements based on the latent diffusion model (LDM) [22]. LiDM first pre-trained an autoencoder to compress the range images and then trained a discrete-time diffusion model on the lower dimensional feature space. RangeLDM [8] is a more recent work following a similar LDM approach. Nevertheless, the LDM approaches still struggle with blurriness caused by the non-iterative decoding by the autoencoder.
In summary, the diffusion model-based approach generates high-fidelity samples with stable training, among the generative model frameworks. However, the sampling process requires a sufficiently large number of steps because the generative process is defined by stochastic curved trajectories defined as SDEs. If these trajectories are approximated with too few steps, the generated LiDAR samples are prone to discretization errors (see Fig. 1 for an example). In this paper, we address this issue by introducing easy-to-approximate trajectories demonstrated in natural image domains [11].
III Method
We employ rectified flow [11] and its extension [12] to construct the generative trajectories that are optimized to be straight so that sampling can be performed in few-step approximation. In this section, we first introduce the procedure to build the straight trajectories by rectified flows and then describe our modifications to LiDAR data generation.
III-A Preliminary
Initial training. Suppose the unknown data distribution $p_{1}$ where the dataset samples $\bm{x}_{1}\sim p_{1}$ are only accessible and Gaussian distribution $p_{0}=\mathcal{N}(0,\bm{I})$ which draws the latent variables $\bm{x}_{0}\sim p_{0}$. Both $p_{1}$ and $p_{0}$ are defined over $\mathbb{R}^{d}$. The goal is to build the transport map between the two distributions. In rectified flows, we consider the following ordinary differential equation (ODE) which transports the samples within the timesteps $t\in[0,1]$:
$d\bm{x}_{t}=v_{\theta}\left(\bm{x}_{t},t\right)dt$ (1)
where $\bm{x}_{t}\in\mathbb{R}^{d}$ is an intermediate state at $t$ and $v_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a neural network to predict the velocity fields towards $\bm{x}_{1}$. To minimize the discretization errors in ODE integration, we consider the linear interpolation path between $\bm{x}_{0}$ and $\bm{x}_{1}$: $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$. Fig. 3(a) illustrates the example trajectory. Then we train the network $v_{\theta}$ by minimizing the following conditional flow matching loss $\mathcal{L}_{\rm{CFM}}$ for independently sampled pairs $(\bm{x}_{1},\bm{x}_{0})$, so that the network $v_{\theta}$ to force the sample $\bm{x}_{t}$ to follow the uniform velocity $\bm{x}_{1}-\bm{x}_{0}$ as much as possible [11].
$\mathcal{L}_{\rm{CFM}}=\mathbb{E}\left[\|\left(\bm{x}_{1}-\bm{x}_{0}\right)-v_{\theta}\left(\bm{x}_{t},t\right)\|^{2}_{2}\right],$ (2)
where $t\sim\mathrm{Uniform}(0,1)$ and $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$. The initial model $v_{\theta}$ obtained from Eq. 2 is referred to as 1-RF in this paper.
Straightening. Although the initial model 1-RF is capable of producing high-quality samples, the built trajectories are not straight because the model is trained on independently sampled pairs $(\bm{x}_{1},\bm{x}_{0})$. As a result, 1-RF still requires a large number of steps for sampling. Rectified flows address this issue by iteratively refining the flow field through reflow [11]. In the reflow process, $\bm{x}_{0}$ is sampled from $p_{0}$, while the target point $\bm{x}_{1}$ is obtained by solving Eq. 1, using 1-RF and the initial value $\bm{x}_{0}$. Training $v_{\theta}$ with the newly dependent pairs $(\bm{x}_{1},\bm{x}_{0})$ reduces the transport cost between $p_{0}$ and $p_{1}$, leading the sample trajectories to become straighter [11]. Following the improved technique proposed by Lee et al. [12], we switch the loss to the following pseudo-Huber loss:
$\mathcal{L}_{\rm{PH}}=\mathbb{E}\left[\sqrt{\|\left(\bm{x}_{1}-\bm{x}_{0}\right)-v_{\theta}\left(\bm{x}_{t},t\right)\|^{2}_{2}+c^{2}}-c\right],$ (3)
where $c=0.00054\sqrt{d}$ and $d$ is the dimension of the data. It is known that the 1-RF loss is difficult to be minimized around $t=0$ and $t=1$ [12]. We will also observe the same phenomenon in Fig. 7. To give more weights for the timesteps, we sample $t$ from the U-shaped distribution [12]: $p_{t}(u)\propto e^{au}+e^{-au}$ where $u\in[0,1]$ and $a=4$. The model obtained by reflow is called 2-RF in this paper.
Timestep distillation. The straightened 2-RF model can be further improved by timestep distillation [11]. At this stage, the model training focuses on the specific timesteps required for few-step sampling, sacrificing predictions at the other unnecessary timesteps. For instance, distilling to a 2-step sampling involves training the model only at $t\in\{0,0.5\}$. All other settings remain the same as in the 2-RF. We denote the $i$-RF model distilled with $k$-step as $i$-RF + $k$-TD.
Sampling. Sampling can be performed by solving the initial value problem of the ODE described in Eq. 1, using the learned velocity estimators $v_{\theta}$. For 1-RF and 2-RF, we can choose the arbitrary number of sampling steps. In general, the larger the number of steps, the smaller the discretization error. For $k$-TD models, the number of sampling steps is fixed at $k$. Any integration solvers can be used for sampling, such as the following Euler methods:
$\bm{x}_{t_{n+1}}\leftarrow\bm{x}_{t_{n}}+\left(t_{n+1}-t_{n}\right)v_{\theta}\left(\bm{x}_{t_{n}},t_{n}\right),$ (4)
where $10\leq t_{n}<t_{n+1}<$ 1 and $\bm{x}_{0}\sim N(0,I)$.
Inversion. The rectified flows can perform inversion, a process of transforming given data $\bm{x}_{1}$ into the corresponding embedding $\bm{x}_{0}$ in the latent space, by solving ODE in Eq. 1 reversely, from $t=1$ to $t=0$. Similar to the other generative models, the inverted $\bm{x}_{0}$ can be used for various applications such as image manipulation. We showcase the application of LiDAR scene interpolation using our model in Fig. 2.
Inversion Spherical linear interpolation (slerp) on latent space Inversion
Figure 2: Scene interpolation using R2Flow inversion. The both side were reconstructed from real samples via inversion. The middle four samples were generated from interpolated latent variables.
III-B Data Representation
Following the existing studies [1, 2, 3, 4, 5, 6], R2Flow is trained on the equirectangular image representation of LiDAR data. We assume a LiDAR sensor that has an angular resolution of $W$ in azimuth and $H$ in elevation and measures the range and reflectance at each laser angle. Then, $HW$ sets of the range and reflectance values can be projected to a 2-channel equirectangular image $\bm{x}_{1}\in\mathbb{R}^{2\times{H}\times{W}}$ by spherical projection. Moreover, following the baselines [5, 4], we also re-scale the range modality $\bm{x}_{\rm{range}}\in[0,x_{\rm{max}}]^{1\times{H}\times{W}}$ into the log-scale representation $\bm{x}_{\rm{log}}\in[0,1]^{1\times{H}\times{W}}$ as follows:
$\bm{x}_{\rm{log}}=\frac{\mathrm{log}(\bm{x}_{\rm{range}}+1)}{\mathrm{log}(x_{\mathrm{max}}+1)}.$ (5)
This log-scale representation gains the geometric resolution of nearby points. Generated range images can be projected back to the 3D point clouds with the reflectance values.
III-C Velocity Estimator
(a) Straight flows (b) HDiT [15]-based overall architecture (c) Details of the building blocks
Figure 3: Schematic overview of our velocity estimator. (a) Straight flows are learned to transport samples between the latent space $p_{0}$ and the image space $p_{1}$. (b) Overall architecture to estimate the velocity fields $\bm{v}_{t}$ from the intermediate state $\bm{x}_{t}$ and the timestep $t$. The Interp layers fuse the current tokens and skipped tokens at each spatial location with learnable weights. (c) The details of the building blocks. The Circular MHSA (multi-head self-attention) layer uses a global attention kernel at the bottleneck and a sliding local window [23] for other stages.
In this section, we describe the design choice of the velocity estimator $v_{\theta}$ in Eq. 1. Fig. 3 depicts the schematic diagram of the model architecture.
Pixel-space vs. feature-space. As discussed in Sec. II, the pixel-space iteration involves high computational cost in general, which barriers the adoption of powerful backbone models such as Vision Transformers (ViTs) [24]. A common approach to this issue is to reduce the dimensionality at which iterative models are trained. In diffusion models, Rombach et al. [22] proposed LDM (latent diffusion model), which consists of an autoencoder (AE) pretrained to perceptually compress images and a diffusion model trained on the lower-dimensional AE features. The compression is motivated by the observation that representing imperceptible details of natural images can be relegated to AE. However, the AE-based non-iterative decoding can be problematic in LiDAR range image generation which requires accurate pixel values and their alignment to maintain geometric fidelity in point clouds. For instance, LiDM [6] proposed the LDM-based approach in the LiDAR generation task but also identified blurry patterns output by the AE decoder as a remaining issue (see Figs. 1 and 6). Therefore, we rethink powerful yet efficient pixel-space architectures for precise modeling and aim to reduce the number of iteration as described in Section III-A.
Architecture design. Our model is built upon HDiT (hourglass diffusion transformer) [15], which is a ViT-based architecture proposed for pixel-space diffusion models. The key idea is to use a sliding window self-attention mechanism [23] to avoid increasing the $\mathcal{O}(n^{2})$ computation with respect to the number of tokens $n$. It thereby enables high-resolution image generation without using the aforementioned compression [22] or staged upsampling [25], despite the pure Transformer structure. In this paper, to generate panoramic range and reflectance images representing LiDAR data, we modify HDiT as follows. (i) The sliding window in the self-attention layers [23] is modified to operate in a horizontal circular pattern using a circular padding technique [5, 4]. (ii) Following the recent ViT-based architecture for LiDAR processing [26], the patch size in tokenization is changed from the default square shape to a landscape shape of $1\times 4$. The sliding windows also has the landscape shape of $3\times 9$. (iii) We use pre-defined LiDAR beam angles to condition the relative positional embeddings (RoPE [27]) in the self-attention layers, limiting the angular frequencies to harmonics. (iv) Similar to ViT [24], we apply a learnable additive bias to the tokens as an absolute positional embedding (APE); otherwise, the generated LiDAR point clouds involve random azimuth rotation.
IV Experiments
In this section, we present the quantitative and qualitative evaluation of the unconditional generation task, focusing on the faithfulness of the sampled LiDAR data.
IV-A Settings
Dataset. Following baselines [5, 6, 4], we utilize the KITTI-360 [16] dataset. The KITTI-360 dataset contains 81,106 point clouds captured using a Velodyne HDL-64E (64-beam mechanical LiDAR sensor). We adopt the standard data split defined by Zyrianov et al. [5]. Each point cloud is projected onto a $64\times 1024$ image, with range and reflectance values assigned to each pixel.
Baselines. We selected baseline methods for which implementations are publicly available. For GAN-based approaches, we compare the vanilla GAN [1] (stable version in [2]), DUSty v1 [2], and DUSty v2 [3]. For diffusion-based approaches, we include comparisons with LiDARGen [5], R2DM [4], and LiDM [6]. We re-trained the GAN models on the KITTI-360 dataset both with and without the reflectance modality. For LiDARGen and R2DM, we used the available pre-trained weights to generate samples. Additionally, we trained LiDM (excluding the autoencoder part) using the same training split, alongside an improved model described in the following.
LiDM improvements. We found that LiDM [6] lacks absolute positional bias in the horizontal direction, leading to random azimuth rotation in the unconditionally generated samples (see Fig. 4 for the visualization). To ensure a fair comparison, we incorporate an absolute positional embedding (APE) into the diffusion model, similar to recent ViT-based diffusion models [28, 29] and ours. As APE, we add learnable biases $\mathbb{R}^{256\times 16\times 128}$ after the first convolution layer and re-train the latent diffusion model. As shown in Fig. 4, the inclusion of APE improved the spatial alignment of the LiDM samples.
Refer to caption Refer to caption Refer to caption
LiDM w/o APE LiDM w/ APE Dataset
Figure 4: Distribution of point clouds generated by LiDM [6]. We calculated a marginal distribution of 1000 random samples. With APE, the distribution get closer to the dataset.
Evaluation metrics. Following the related work, we evaluate the distributional similarity between real and generated samples across multiple levels of data representation. We use seven evaluation metrics: Fréchet range distance (FRD)[5], Fréchet range image distance (FRID)[6], Fréchet point cloud distance (FPD)[30], Fréchet point-based volume distance (FPVD)[6], Fréchet sparse volume distance (FSVD)[6], Jensen–Shannon divergence (JSD)[5], and maximum mean discrepancy (MMD) [5], based on five types of data representations: range images, reflectance images, point clouds, voxels, and bird’s-eye views (BEV). The point cloud, voxel, and bird’s-eye view representations are derived from the range image. Table I summarizes the correspondence between the metrics and the representations. Note that only FRD incorporates both range and reflectance modalities, while the other metrics rely solely on the range modality. For each method, we generate 10,000 samples and evaluate them against the entire dataset of real samples, following standard practices in generative models.
Implementation details. All models are implemented using PyTorch. We performed a distributed training with automatic mixed precision (AMP). We used torchdiffeq [31] for solving ODEs. For training 2-RF, we sampled 1M pairs by the dopri5 sampler (adaptive step-size) with absolute/relative tolerance of 1e-5. For timestep distillation, we sampled 100k pairs with the same sampler. All evaluation results are produced by the euler sampler (fixed step-size) for fair comparison. Our code and pretrained weights will be available on https://github.com/kazuto1011/r2flow.
IV-B Results
Quantitative results. Tab. I shows the evaluation results in three groups from top to bottom In the first group which compares GAN-based methods, DUSty v2 that generates both range and reflectance images performed well across multiple metrics. The second group compares the results of the iterative models including ours, with higher NFEs (number of function evaluations). NFE counts the number of times running neural networks. With the introduction of APE, LiDM demonstrates improvements across all metrics, particularly in BEV-based JSD and MMD. Our R2Flow achieved results comparable to another pixel-space model, R2DM. On the other hand, the 2-RF scores were slightly lower, which is because the quality upper bound for 2-RF is limited by the parent model 1-RF rather than real data. Overall, the iterative models are better than the GANs. The third group compares the results with fewer NFEs. While all baseline methods exhibit significant performance degradation, our R2Flow mitigates the degradation through reflow and distillation. By incorporating the few-step distillation (2-TD and 4-TD), some metrics exhibit results comparable to those achieved with a larger number of steps. Fig. 5 shows FRD scores as a function of NFEs in details. R2Flow shows the better computational tradeoff.
TABLE I: Quantitative Evaluation of Unconditional Generation on KITTI-360
Method Output Framework NFE $\downarrow$ FRD $\downarrow$ FRID $\downarrow$ FPD $\downarrow$ FPVD $\downarrow$ FSVD $\downarrow$ JSD $\downarrow$ MMD $\downarrow$
Vanilla GAN [1] GAN 0001 - 221.89 1258.86 268.16 358.28 10.09 007.22
GAN 0001 2766.43 216.45 1367.85 196.35 257.98 08.73 005.22
DUSty v1 [2] GAN 0001 - 174.74 0102.11 100.91 118.23 03.90 002.19
GAN 0001 1429.96 194.25 0318.06 134.24 171.41 05.40 002.95
DUSty v2 [3] GAN 0001 - 140.23 0713.59 096.69 120.00 05.17 003.22
GAN 0001 1132.02 180.71 0641.12 088.72 106.34 02.98 000.87
LiDARGen [5] SMLD 1160 0542.74 066.45 0090.29 043.56 050.14 04.04 001.74
LiDM (original) [6] DDPM 0050 - 009.21 0503.74 019.45 018.08 17.40 033.15
LiDM [6] + APE DDPM 0050 - 006.21 0393.18 015.25 015.87 04.50 000.89
DDPM 0200 - 005.75 0372.81 014.65 015.20 04.30 000.67
R2DM [4] DDPM 0256 0154.11 003.70 0003.79 010.90 009.11 02.19 000.73
R2Flow (ours) 1-RF 0256 0122.81 004.18 0009.32 016.51 015.07 02.16 000.30
2-RF 0256 0148.09 008.64 0011.06 029.08 029.20 02.24 000.29
LiDM [6] + APE DDPM 0001 - 191.00 1240.25 138.43 200.10 36.14 188.41
R2DM [4] DDPM 0001 2981.89 237.41 0237.40 378.92 458.69 33.48 089.14
R2Flow (ours) 1-RF 0001 2724.67 254.03 3967.51 374.89 398.95 33.76 096.90
2-RF 0001 0743.97 119.42 0027.94 090.92 107.72 03.86 001.11
2-RF + 1-TD 0001 0336.62 054.75 0013.08 064.03 070.15 02.37 000.35
2-RF + 2-TD 0002 0212.08 022.57 0011.00 045.56 048.98 02.39 000.31
2-RF + 4-TD 0004 0187.10 015.92 0010.92 038.64 040.75 02.40 000.31
  • Notation: range image, reflectance image, point cloud, voxel, bird’s eye view.
  • For each group, we highlight the top-1 scores in bold and the top-2 scores in shaded. The JSD and MMD scores are multiplied by $10^{2}$ and $10^{4}$, respectively.
Figure 5: Speed–quality tradeoff in FRD. We compare the methods [2, 3, 5, 4] supporting both range and reflectance modalities. Our R2Flow (blue lines) shows the better tradeoff against the baselines (black lines).
Qualitative results. In Fig. 6, we compare generated LiDAR point clouds from real data [16], DUSty v2 [3], LiDARGen [5], our improved LiDM [6], R2DM [4], and our R2Flow. LiDM, R2DM, and R2Flow demonstrate better quality, such as sharper scan lines and clearer object boundaries, while LiDM exhibits some wavy and blurry boundaries, as previously reported by Ran et al. [6].
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Training data DUSty v2 [3] LiDARGen [5] LiDM [6] + APE R2DM [4] R2Flow (ours)
(KITTI-360 [16]) (GAN, NFE = 1) (SMLD, NFE = 1160) (DDPM, NFE = 200) (DDPM, NFE = 256) (1-RF, NFE = 256)
Figure 6: Comparison of unconditional generation results.
Model architecture. In Table II, we compare three model architectures for the pixel-space velocity estimator: Efficient U-Net [25] (CNN) as used in R2DM [4], ADM U-Net [32] (CNN) commonly used for natural images, and HDiT [15] (Transformer) as used in ours. Among the tested configurations, the HDiT-based architecture achieved the best performance. Although increasing the number of parameters improves the performance of both Efficient U-Net and ADM U-Net, bringing them closer to HDiT, this comes at the cost of higher computational complexity and increased latency.
TABLE II: Architecture Comparison of Velocity Estimator
Base architecture FLOPs (G) Params (M) Latency (ms) FRD [5]
Efficient U-Net [25] 116.3 031.1 15.2 151.90
+ larger model size 688.3 284.6 39.5 124.49
ADM U-Net [32] 265.6 087.4 24.8 140.66
+ larger model size 692.7 125.5 50.2 134.22
HDiT [15] 077.8 080.9 28.8 122.81
  • We trained the 1-rectified flow with the different architectures and evaluated FRD with the 256-step Euler sampling.
Trajectory curvature. To verify the straightening effect by reflow, we measure the trajectory curvature over timestep, defined in prior work [11, 12]:
$s\left(t\right)=\|\left(\Phi\left(\bm{x}_{0},1\right)-\bm{x}_{0}\right)-v_{\theta}\left(\Phi\left(\bm{x}_{0},t\right),t\right)\|^{2}_{2},$ (6)
where $\bm{x}_{0}\sim p_{0}$ and $\Phi\left(\bm{x}_{t_{\mathrm{start}}},t_{\mathrm{end}}\right)$ is the ODE solution from the timestep $t_{\mathrm{start}}$ to $t_{\mathrm{end}}$ with the initial value $\bm{x}_{t_{\mathrm{start}}}$. As the trajectory becomes straighter, $s(t)$ approaches zero. Fig.7(a) illustrates the trajectory curvature for 1-RF and 2-RF over 256 timesteps. 1-RF exhibits high curvature at both early and late timesteps. In Fig. 7(b), we visualize the top 200 most curved trajectories. It is evident that many 1-RF trajectories exhibit significant curvature, necessitating numerous sampling steps for generating high-quality samples. Notably, a part of the trajectories with high curvature near $-1$ correspond to pixels affected by raydrop noise. The one-time reflow in 2-RF markedly enhances trajectory straightness, as also validated by quantitative results in Tab. I.
(a) Curvature over time (b) Trajectories of pixels
Figure 7: Trajectory curvature of learned flows. (a) Trajectory curvatures between 1-RF and 2-RF (b) top-200 curved trajectories (0.15% of all pixels). The pixel value $-1$ at $t=1$ corresponds to raydrop noise.
V Conclusions
In this paper, we presented R2Flow, the rectified flow-based generative model for fast and realistic LiDAR data generation. We verified the effectiveness of our approach in both efficiency and quality in the unconditional generation evaluation. Future work will focus on investigating the scalability of R2Flow, improving the reflow process to maintain quality, and demonstrating R2Flow in application tasks such as sparse-to-dense completion, sim-to-real domain adaptation, and anomaly detection. The trajectory visualization in Fig. 7 suggests that the raydrop pixels drifting toward a value of $-1$ may hinder training straight flows. We expect that implementing a raydrop-aware architecture [2, 3] could mitigate this issue.